# Lean Copilot: Large Language Models as Copilots for Theorem Proving in Lean

### v2

@@ Abstract
Neural theorem proving combines large language models (LLMs) with proof assistants such as Lean, where the correctness of formal proofs can be rigorously verified, leaving no room for hallucination. With existing neural theorem provers pretrained on a fixed collection of data and offering valuable suggestions at times, it is challenging for them to continually prove novel theorems in a fully autonomous mode, where human insights may be critical. In this paper, we explore LLMs as copilots that assist humans in proving theorems. We introduce Lean Copilot, an general framework for running LLM inference natively in Lean. It enables programmers to build various LLM-based proof automation tools that integrate seamlessly into the workflow of Lean users. Lean users can use our pretrained models or bring their own ones that run either locally (with or without GPUs) or on the cloud. Using Lean Copilot, we build LLM-based tools that suggest proof steps, complete proof goals, and select relevant premises. Experimental results on the Mathematics in Lean textbook demonstrate the effectiveness of our method compared to existing rule-based proof automation in Lean (aesop), confirming the significance of having LLMs integrated into the theorem proving workflow in Lean. When assisting humans, Lean Copilot requires only 2.08 manually-entered proof steps on average (3.86 required by aesop); when automating the theorem proving process, Lean Copilot automates 74.2% proof steps on average, 85% better than aesop (40.1%). We open source all code and artifacts under a permissive MIT license to facilitate further research.
@@
Neural Theorem Proving, Proof Automation, Large Language Models, Neuro-Symbolic Reasoning
@@ 1 Introduction
Lean Copilot with our automation tools can be installed easily as a Lean package and run on most hardware, paving the way for wide adoption and impact on the Lean community (Appendix D). They are among the first steps in making LLMs accessible to human users of proof assistants, which we hope will initiate a positive feedback loop where proof automation leads to better data and ultimately improves LLMs on mathema. We open source all code on Github (Appendix A) for Lean Copilot at https://github.com/lean-dojo/LeanCopilot to facilitate future research. We make all evaluation code and detailed results public at https://github.com/Peiyang-Song/mathematics_in_lean/tree/full-scale-experiment for full transparency.
@@ 4.1 Generating Tactic Suggestions
When humans prove theorems in Lean, they inspect the remaining proof goals to decide for the next tactic (i.e. proof step in Lean). Tactics do not come from a predefined list; they are more like programs in a domain-specific language (DSL). They can take arbitrary Lean terms as parameters, and simpler tactics can be combined into compound ones. Users can also extend existing tactics by defining customized tactics. Due to these complexities, producing the right tactic can be challenging even for experienced Lean users.
@@ 6 Conclusion
Peiyang Song was supported by the Summer Undergraduate Research Fellowship at Caltech. Kaiyu Yang was supported by the Computing, Data, and Society Postdoctoral Fellowship at Caltech. Anima Anandkumar is partially supported by the Bren endowed chair and Schmidt Sciences AI2050 senior fellowship. We appreciate the valuable feedback from members of the Anima AI+Science Lab on an initial version of this paper, and from the Lean community on the releases of our tool. We thank Scott Morrison for suggestions on simplifying Lean Copilot’s installation and Mac Malone for helping implement it. We thank Jannis Limperg for supporting our LLM-generated tactics in Aesop.
@@ Appendix B Generalizing Lean Copilot with Low-Level Interfaces
To run LLMs in Lean through server processes, users can declare an ExternalGenerator, where they specify a model to use, a host name, and a port number. Then once they open the corresponding server in the command line, the specified model will be able to perform text-to-text generation, and display the outputs on the infoview just like the native generators do in Fig. 5. Text-to-vector encoding can be performed in the same way, by declaring an ExternalEncoder. In principle, users can bring any models using Lean Copilot through ExternalGenerator and ExternalEncoder. In order to use them, users just need to expose certain APIs such as model name, input sequence, generation prefix, etc. We provide examples of using this Python API server in our codebase, which current include running models from OpenAI’s GPT family (OpenAI, 2024), Anthropic’s Claude family (Anthropic, 2024), Google’s Gemini family (Team, 2024), and InternLM2 (Cai et al., 2024).

### v3

@@ Abstract
Neural theorem proving combines large language models (LLMs) with proof assistants such as Lean, where the correctness of formal proofs can be rigorously verified, leaving no room for hallucination. With existing neural theorem provers pretrained on a fixed collection of data and offering valuable suggestions at times, it is challenging for them to continually prove novel theorems in a fully autonomous mode, where human insights may be critical. In this paper, we explore LLMs as copilots that assist humans in proving theorems. We introduce Lean Copilot, a general framework for running LLM inference natively in Lean. It enables programmers to build various LLM-based proof automation tools that integrate seamlessly into the workflow of Lean users. Lean users can use our pretrained models or bring their own ones that run either locally (with or without GPUs) or on the cloud. Using Lean Copilot, we build LLM-based tools that suggest proof steps, complete proof goals, and select relevant premises. Experimental results on the Mathematics in Lean textbook demonstrate the effectiveness of our method compared to existing rule-based proof automation in Lean (aesop), confirming the significance of having LLMs integrated into the theorem proving workflow in Lean. When assisting humans, Lean Copilot requires only 2.08 manually-entered proof steps on average (3.86 required by aesop); when automating the theorem proving process, Lean Copilot automates 74.2% proof steps on average, 85% better than aesop (40.1%). We open source all code and artifacts under a permissive MIT license to facilitate further research.
@@
Neural Theorem Proving, Proof Automation, Large Language Models, Neuro-Symbolic Reasoning, AI for Mathematics
@@ 1 Introduction
Lean Copilot with our automation tools can be installed easily as a Lean package and run on most hardware, paving the way for wide adoption and impact on the Lean community (Appendix D). They are among the first steps in making LLMs accessible to human users of proof assistants, which we hope will initiate a positive feedback loop where proof automation leads to better data and ultimately improves LLMs on mathema. We open source all code on Github (Appendix A) at https://github.com/lean-dojo/LeanCopilot to facilitate future research. We make all evaluation code and detailed results public at https://github.com/Peiyang-Song/mathematics_in_lean/tree/full-scale-experiment for full transparency.
@@ 4.1 Generating Tactic Suggestions
When humans prove theorems in Lean, they inspect the remaining proof goals to decide for the next tactic (i.e. proof step in Lean). Tactics do not come from a predefined list; they are similar to programs in a domain-specific language (DSL). They can take arbitrary Lean terms (or take none) as parameters, and simpler tactics can be combined into compound ones. Users can also extend existing tactics by defining customized tactics. Due to these complexities, producing the right tactic can be challenging even for experienced Lean users.
@@ 6 Conclusion
Peiyang Song was supported by the Summer Undergraduate Research Fellowship at Caltech. Kaiyu Yang was supported by the Computing, Data, and Society Postdoctoral Fellowship at Caltech. Anima Anandkumar is partially supported by the Bren endowed chair and Schmidt Sciences AI2050 senior fellowship. We appreciate the valuable feedback from members of the Anima AI+Science Lab on an initial version of this paper, from the Lean community on the releases of our tool, and from the anonymous reviewers at NeuS 2025. We thank Scott Morrison for suggestions on simplifying Lean Copilot’s installation and Mac Malone for helping implement it. We thank Jannis Limperg for supporting our LLM-generated tactics in Aesop.
@@ Appendix B Generalizing Lean Copilot with Low-Level Interfaces
To run LLMs in Lean through server processes, users can declare an ExternalGenerator, where they specify a model to use, a host name, and a port number. Then once they open the corresponding server in the command line, the specified model will be able to perform text-to-text generation, and display the outputs on the infoview just like the native generators do in Fig. 5. Text-to-vector encoding can be performed in the same way, by declaring an ExternalEncoder. In principle, users can bring any models using Lean Copilot through ExternalGenerator and ExternalEncoder. In order to use them, users just need to expose certain APIs such as model name, input sequence, generation prefix, etc. We provide examples of using this Python API server in our codebase, which current include running models from OpenAI’s GPT family (OpenAI, 2024), Anthropic’s Claude family (Anthropic, 2024), Google’s Gemini family (Team, 2024), InternLM2 (Cai et al., 2024), etc.